D | Question 26 Refer to the following partial ANOVA results from Excel (some information is missing). Source of Variation SS MS 210.2778 74 15 df etween groups Within groups Total 1483 2113 833 The number of treatment groups is: O 2. 3. Previous
Answer #1

SSbetween=SStotal-SSwithin = 2113.833-1483=630.833

MSbetween =SSbetween/df

df=SSbetween/MSbetween = 630.833/210.2778 =2.999999

so, Df~3

so, no of treatments = 3+1 = 4